# Math Reasoning in Solid Geometry

### v1

@@ 2 Related Works
With the advancement of LLMs and vision-language alignment techniques, early multimodal models such as MiniGPT-4 minigpt , LLaMA-Adapter llamaadapter , and LLaVA llava demonstrated promising capabilities in visual understanding. More recent models such as closed-source GPT-4o GPT4o , Claude claude , Gemini gemini and open-source Qwen-VL qwen2d5vl , InternVL internvl , and LLaVA-Onevision llavaonevision have further pushed the boundaries of general-purpose visual reasoning. The recent release of OpenAI’s o1 model has highlighted the effectiveness of long Chain-of-Thought in improving reasoning performance, this has inspired a wave of O1-style models  virgo ; visualrft ; VLRethinker ; fast ; r1vl ; reasonrft ; Openvlthinker , many of which show strong performance on multimodal mathematical tasks.
@@ 3.2 Data Construction
GeoQA CN 0 0.0% 100.0% KK SS
Geometry3K EN 0 0.0% 100.0% KK SS
UniGeo EN 0 0.0% 100.0% KK SS
PGPS9K EN 0 0.0% 100.0% KK SS
MMMU-MATH EN 0 0.0% 29.3% UU SS
GeoEval EN 100 2.0% 94.1% KK PP
DynaMath EN 150 3.0% 15.4% KK SS
MATH-Vision EN 263 8.7% 58.7% KK UU SS
OlympiadBench CN 784 9.2% 15.6% CC SS
MathVerse EN 119 15.1% 64.7% KK SS PP
GeoSense EN/CN 350 20.0% 80.0% KK SS PP 2
@@ 3.3 Comparison with Existing Benchmarks
In contrast, SolidGeo is the first large-scale benchmark dedicated to solid geometry. Each problem is paired with at least one visual input and has been manually verified for correctness. Unlike prior datasets that use broad or ambiguous labels, SolidGeo introduces a refined taxonomy with eight fine-grained categories, explicitly capturing the diversity of solid geometry subdomains. Furthermore, SolidGeo features a significantly longer average question length (77.2 tokens), compared to MathVista (15.6) and MathVision (42.3), indicating higher contextual richness and greater reasoning complexity. Together, these attributes make SolidGeo a comprehensive and challenging benchmark for evaluating multimodal mathematical reasoning in 3D spatial contexts. A detailed comparison with existing benchmarks is presented in Table 2. See Appendix C for details.
@@ 4.2 Experimental Results
Failure of CoT/few-shot. As shown in Table 4, Chain-of-Thought (CoT) and few-shot prompting do not consistently improve model performance on SolidGeo. Models like Qwen2.5-VL-72B and Gemini-1.5-pro benefit from CoT and 2-shot settings, with accuracy increasing from 24.2% to 28.8% and 26.1%, and from 25.3% to 26.1% and 27.5%, respectively. However, other models such as GPT-4o and LLaVA-OV-72B exhibit performance degradation under the same conditions. To further understand this, we analyze the effect of prompting strategies on the average number of generated tokens, as shown in Figure 4. For many models, neither CoT nor 2-shot leads to longer outputs, suggesting such reasoning capabilities may already be learned during pretraining. Notably, we observe that only when CoT increases token count by over 10% does it yield tangible gains—e.g., Qwen2.5-VL-72B: tokens ↑21.9%, accuracy ↑4.6%; LLaVA-OV-7B: tokens ↑20.6%, accuracy ↑1.9%. In contrast, 2-shot prompting shows no consistent correlation with performance improvement, indicating it is not a universally effective strategy.
@@ 5 Conclusion
Limitations and Future Works. SolidGeo has certain limitations. First, the dataset includes only English and Chinese problems, which may limit its applicability for evaluating multilingual reasoning capabilities. Second, there currently exists no widely adopted formal language to represent solid geometry problems, making it difficult to standardize symbolic reasoning in this domain. Future work could explore designing a formal representation framework for solid geometry, which may further facilitate precise modeling and programmatic evaluation.
@@ Appendix C Comparison
As shown in Figure 12, our SolidGeo exhibits the longest average question length among all compared benchmarks, reaching 77.2 tokens. This is significantly higher than datasets such as MathVista (15.6), GeoQA (37.1), and MMMU-MATH (40.8). The increased question length in SolidGeo directly reflects the need for more precise problem understanding and more complex reasoning processes, which distinguishes it from benchmarks that primarily involve shorter and more formulaic tasks. Moreover, the right subfigure highlights a striking performance gap: despite the state-of-the-art models achieving over 80% accuracy on other benchmarks, their best performance on SolidGeo drops sharply to 49.5%. This further validates the higher reasoning difficulty and challenge posed by our benchmark. It underscores the necessity of more robust spatial understanding and compositional reasoning to succeed on SolidGeo.
@@ D.1 Data Collection Details
To construct SolidGeo, we adopted a hybrid data collection strategy that combines both existing benchmarks and newly collected real-world data. Specifically, we extracted solid geometry problems from six publicly available multimodal math datasets: MathVista, MATH-Vision, DynaMath, GeoEval, CMMaTH, and OlympiadBench, based on their official topic annotations or keyword filtering. These samples account for 1,737 problems in our final dataset.

### v2

@@ 2 Related Works
With the advancement of LLMs and vision-language alignment techniques, early multimodal models such as MiniGPT-4 minigpt , LLaMA-Adapter llamaadapter , and LLaVA llava demonstrated promising capabilities in visual understanding. Recent research also explores parameter-efficient fine-tuning (PEFT) approaches such as LoRA lora and its improved orthogonal reinitialization strategies UORA UORA . More recent models such as closed-source GPT-4o GPT4o , Claude claude , Gemini gemini and open-source Qwen-VL qwen2d5vl , InternVL internvl , and LLaVA-Onevision llavaonevision have further pushed the boundaries of general-purpose visual reasoning. The recent release of OpenAI’s o1 model has highlighted the effectiveness of long Chain-of-Thought in improving reasoning performance, this has inspired a wave of o1-style models  virgo ; visualrft ; VLRethinker ; fast ; r1vl ; reasonrft ; Openvlthinker , many of which show strong performance on multimodal mathematical tasks. Beyond step-wise reasoning, Chain-of-Reasoning Chain-of-Reasoning advocates for a unified multi-paradigm approach, integrating symbolic, visual, and CoT-based reasoning for mathematical problem solving.
@@ 3.2 Data Construction
GeoQA geoqa CN 0 0.0% 100.0% KK SS
Geometry3K Geometry3k EN 0 0.0% 100.0% KK SS
UniGeo unigeo EN 0 0.0% 100.0% KK SS
PGPS9K PGPS9K EN 0 0.0% 100.0% KK SS
MMMU-MATH mmmu EN 0 0.0% 29.3% UU SS
GeoEval geoeval EN 100 2.0% 94.1% KK PP
DynaMath dynamath EN 150 3.0% 15.4% KK SS
MATH-Vision mathvision EN 263 8.7% 58.7% KK UU SS
OlympiadBench olympiadbench EN/CN 784 9.2% 15.6% CC SS
MathVerse mathverse EN 119 15.1% 64.7% KK SS PP
GeoSense geosense EN/CN 350 20.0% 80.0% KK SS PP 2
@@ 3.3 Comparison with Existing Benchmarks
In contrast, SolidGeo is the first large-scale benchmark dedicated to solid geometry. Each problem is paired with at least one visual input and has been manually verified for correctness. Unlike prior datasets that use broad or ambiguous labels, SolidGeo introduces a refined taxonomy with eight fine-grained categories, explicitly capturing the diversity of solid geometry subdomains. Furthermore, SolidGeo features a significantly longer average question length (77.2 words), compared to MathVista (15.6) and MathVision (42.3), indicating higher contextual richness and greater reasoning complexity. Together, these attributes make SolidGeo a comprehensive and challenging benchmark for evaluating multimodal mathematical reasoning in 3D spatial contexts. A detailed comparison with existing benchmarks is presented in Table 2. See Appendix C for details.
@@ 4.2 Experimental Results
Failure of CoT/few-shot. As shown in Table 4, Chain-of-Thought (CoT) and few-shot prompting do not consistently improve model performance on SolidGeo. Models like Qwen2.5-VL-72B and Gemini-1.5-pro benefit from CoT and 2-shot settings, with accuracy increasing from 24.2% to 28.8% and 26.1%, and from 25.3% to 26.1% and 27.5%, respectively. However, other models such as GPT-4o and LLaVA-OV-72B exhibit performance degradation under the same conditions. To further understand this, we analyze the effect of prompting strategies on the average number of generated tokens, as shown in Figure 4. For many models, neither CoT nor 2-shot leads to longer outputs, suggesting such reasoning capabilities may already be learned during pretraining. Notably, we observe that only when CoT increases token count by over 10% does it yield tangible gains—e.g., Qwen2.5-VL-72B: tokens ↑21.9%, accuracy ↑4.6%; LLaVA-OV-7B: tokens ↑20.6%, accuracy ↑1.9%. In contrast, 2-shot prompting shows no consistent correlation with performance improvement, indicating it is not a universally effective strategy. Efficient long-context modeling remains a challenge, despite recent advances in scalable context generalization techniques LongRecipe .
@@ 5 Conclusion
Limitations and Future Works. SolidGeo has certain limitations. First, the dataset includes only English and Chinese problems, which may limit its applicability for evaluating multilingual reasoning capabilities zhao2024large . Second, there currently exists no widely adopted formal language to represent solid geometry problems, making it difficult to standardize symbolic reasoning in this domain. Future work could explore designing a formal representation framework for solid geometry, which may further facilitate precise modeling and programmatic evaluation.
@@ Appendix C Comparison
As shown in Figure 12, our SolidGeo exhibits the longest average question length among all compared benchmarks, reaching 77.2 words. This is significantly higher than datasets such as MathVista (15.6), GeoQA (37.1), and MMMU-MATH (40.8). The increased question length in SolidGeo directly reflects the need for more precise problem understanding and more complex reasoning processes, which distinguishes it from benchmarks that primarily involve shorter and more formulaic tasks. Moreover, the right subfigure highlights a striking performance gap: despite the state-of-the-art models achieving over 80% accuracy on other benchmarks, their best performance on SolidGeo drops sharply to 49.5%. This further validates the higher reasoning difficulty and challenge posed by our benchmark. It underscores the necessity of more robust spatial understanding and compositional reasoning to succeed on SolidGeo.
@@ D.1 Data Collection Details
To construct SolidGeo, we adopted a hybrid data collection strategy that combines both existing benchmarks and newly collected real-world data. Specifically, we extracted solid geometry problems from six publicly available multimodal math datasets: MathVerse mathverse , MATH-Vision mathvision , DynaMath dynamath , GeoEval geoeval , CMMaTH cmmath , and OlympiadBench olympiadbench , based on their official topic annotations or keyword filtering. These samples account for 1,737 problems in our final dataset.